OpenStudy (anonymous):

An object projected vertically upward from ground level returns to earth in 22 seconds. Find the initial velocity in feet per second.

OpenStudy (stormfire1):

You're given a time of 22s so you know that half of that will be the exact point at which the object begins to fall back to Earth. This is also the point where the instantaneous velocity is 0. The acceleration throughout the travel will just be g since there are no other forces involved. So start by writing out what you know: a = -9.8 m/s^2 Vf = 0 m/s (@ the top of the object's travel) t = 11 s From your kinematic equations: Vf = Vi + at 0 = Vi + (-9.8 m/s^2)(11 s) Vi = 107.8 m/s

OpenStudy (anonymous):

wait i have anwer choices adn none of theses are the answer

OpenStudy (stormfire1):

hmmm...what are the choices?

OpenStudy (stormfire1):

OH WAIT....I did m/s not ft/s

OpenStudy (stormfire1):

One sec, while I redo it :)

OpenStudy (anonymous):

no probblem...here are teh choices a) 352, b) 368, c) 348, d) 362, e) 358

OpenStudy (stormfire1):

a = -32 ft/s^2 Vf = 0 m/s (@ the top of the object's travel) t = 11 s From your kinematic equations: Vf = Vi + at 0 = Vi + (-32 ft/s^2)(11 s) Vi = 352 ft/s

OpenStudy (stormfire1):

My bad :)

OpenStudy (stormfire1):

Vf = 0 ft/s and not m/s...forgot to correct that. Anyway, hopefully you understand what I did but if not, just ask

OpenStudy (anonymous):

wait may u help me w/ other ones....

OpenStudy (stormfire1):

Maybe but you should post them separately

OpenStudy (anonymous):

ok